# Robust LLM Alignment via Distributionally Robust Direct Preference Optimization

### v2

@@ Abstract
A major challenge in aligning large language models (LLMs) with human preferences is the issue of distribution shift. LLM alignment algorithms rely on static preference datasets, assuming that they accurately represent real-world user preferences. However, user preferences vary significantly across geographical regions, demographics, linguistic patterns, and evolving cultural trends. This preference distribution shift leads to catastrophic alignment failures in many real-world applications. We address this problem using the principled framework of distributionally robust optimization, and develop two novel distributionally robust direct preference optimization (DPO) algorithms, namely, Wasserstein DPO (WDPO) and Kullback-Leibler DPO (KLDPO). We characterize the sample complexity of learning the optimal policy parameters for WDPO and KLDPO. Moreover, we propose scalable gradient descent-style learning algorithms by developing suitable approximations for the challenging minimax loss functions of WDPO and KLDPO. Our empirical experiments using benchmark data sets and LLMs demonstrate the superior performance of WDPO and KLDPO in substantially improving the alignment when there is a preference distribution shift.
@@ 1 Introduction
The alignment of large language models (LLMs) with human values and preferences is a central objective in machine learning, enabling these models to produce outputs that are useful, safe, and aligned with human intent. Since LLMs are trained on vast, diverse datasets using self-supervised learning, an additional alignment phase is often required to refine their behavior based on human feedback. A widely adopted approach for this is Reinforcement Learning from Human Feedback (RLHF) (Christiano et al.,, 2017; Ziegler et al.,, 2019; Ouyang et al.,, 2022), which involves training a reward model using human preference data and optimizing the LLM using reinforcement learning (RL) approaches, such as proximal policy optimization. More recently, Direct Preference Optimization (DPO) has emerged as an alternative that simplifies the alignment process by directly optimizing model parameters based on human preferences without requiring an explicit reward model. These alignment techniques have played a crucial role in improving the ability of LLMs to generate responses that adhere to human expectations and societal norms, leading to today’s powerful chat models (Achiam et al.,, 2023; Touvron et al.,, 2023).
@@
Despite the importance of the LLM alignment problem, RLHF and DPO remain fundamentally challenging and fragile, mainly due to three reasons. $(i)$ Diversity of human preferences: Standard RLHF/DPO approaches implicitly assume that human preferences can be accurately captured by a single reward function. In reality, human preferences are highly diverse, context-dependent, and distributional, making it infeasible to represent them with a one-size-fits-all optimization framework (Zhao et al.,, 2024; Durmus et al.,, 2023). Standard preference-learning methods tend to skew toward the preferences represented in the majority of training data, disproportionately penalizing minority opinions and reinforcing biases (Chakraborty et al.,, 2024). $(ii)$ Reward hacking: The quality of human preference feedback is inherently noisy, ambiguous, and inconsistent, as they are collected from human annotators who may lack domain expertise, exhibit labeling fatigue, or hold conflicting opinions (Zhang et al.,, 2024; Wu et al., 2024a, ), which can often lead to misaligned preference estimation. This issue is exacerbated by reward hacking, where models learn undesirable shortcuts to maximize the estimated reward function, generating responses that appear aligned but deviate from genuine human intent (Amodei et al.,, 2016; Skalse et al.,, 2022; Eisenstein et al.,, 2023). $(iii)$ Distribution shift: Alignment algorithms use static preference datasets for training, collected under controlled conditions. However, the preferences of real-world users can often be out-of-distribution from that of the training data, depending on the geographical region, demography, linguistic patterns, and emerging social trends, among many others. A model aligned using a specific fixed dataset may fail catastrophically when deployed to users whose preference distribution does not match that of the training data (Casper et al.,, 2023; LeVine et al.,, 2023; Kirk et al.,, 2024).
@@
In this paper, we address the fragility of the LLM alignment using DPO, with a particular focus on the challenges arising from the preference distribution shift. DPO reduces the alignment problem to a supervised learning problem. It is known that the performance of supervised learning algorithms degrades significantly in the out-of-distribution setting (Taori et al.,, 2020; Koh et al.,, 2021), which is exacerbated due to the realistic distribution shift scenarios arising in the LLM deployment. Distributionally robust optimization/learning framework has been recently used to address the issue of distribution shift in various settings (Duchi and Namkoong,, 2021; Kuhn et al.,, 2019; Chen et al.,, 2020). This framework considers an uncertainty set of data distributions around a nominal distribution (typically the training data distribution) and solves a minimax optimization problem to minimize the expected loss, where the expectation is taken with respect to the distribution in the uncertainty set that maximizes the loss. The distributionally robust learning approach has been successfully applied, with theoretical guarantees and scalable algorithms, in supervised learning (Chen and Paschalidis,, 2018; Namkoong and Duchi,, 2016; Levy et al.,, 2020), multi-armed bandits (Si et al.,, 2020; Yang et al.,, 2023), and reinforcement learning (Wang and Zou,, 2022; Panaganti et al.,, 2022; Zhou et al.,, 2024; Xu et al.,, 2023). This motivates us to address the following questions:
@@
    Empirically, we show that standard DPO is sensitive to preference distribution shift, leading to degraded performance when training and evaluation rewards differ. In contrast, our robust variants, WDPO and KLDPO, consistently achieve superior performance across diverse alignment tasks. For example, we fine-tune LLaMA-3.2-1B/3B-Instruct and LLaMA-3.1-8B-Instruct models on prompts from the HelpSteer2 dataset (Wang et al., 2024b, ) using preferences generated by the ArmoRM reward model (Wang et al., 2024a, ), and evaluate them on distinct reward objectives from the OpenLLM Leaderboard (Fourrier et al.,, 2024).
@@ 2 Related Work
Robust RLHF: Bai et al., (2022) proposed to adjust weights on the combination of loss functions based on different topics (harmless vs. helpful) for robust reward learning. Chakraborty et al., (2024) proposed to learn multiple reward functions for different sub-populations through an expectation-maximization approach, and a robust policy based on these rewards via a max-min optimization, which is different from our distributional robust learning approach. Padmakumar et al., (2024) augmented the existing binary preference datasets with synthetic preference judgments to estimate the diversity of user preferences. Yan et al., (2024) proposed a Bayesian reward model ensemble to quantify the uncertainty of reward estimation and used it to reduce reward overoptimization. Bukharin et al., (2024) proposed a robust RLHF approach for addressing the preference data corruption problem.
Robust DPO: Huang et al., (2024) proposed $\chi$PO that implements the principle of pessimism in the face of uncertainty via regularization with the $\chi^{2}$-divergence for avoiding reward hacking/overoptimization w.r.t. the estimated reward. Ramesh et al., (2024) proposed Group Robust Preference Optimization (GRPO) to address the diverse preference problem by modeling the total loss as a weighted sum of individual DPO losses computed on separate preference datasets, and optimizing for the worst-case weighting. In contrast, our approach does not assume access to such a group structure and instead directly models distributional robustness over a single dataset that implicitly aggregates diverse preferences. Chowdhury et al., (2024) considered the setting where $\epsilon$-fraction of the preference labels in the training dataset is corrupted and proposed a noise-robust algorithm to mitigate its effect, assuming the knowledge of $\epsilon$. Wu et al., 2024b focused on adapting the DPO penalty parameter $\beta$ to handle varying data quality within the training set. The most related work is Wu et al., 2024a , which applies distributional robustness to mitigate data corruption and noise in preference data. Unlike our work, it does not address distribution shift or provide theoretical guarantees, and lacks empirical evaluation on preference distribution shift. Concurrent to our work, Mandal et al., (2025) proposed a distributionally robust version of RLHF and DPO, using total variation (TV) uncertainty sets. However, their theoretical analysis offers the natural policy gradient (NPG) style optimization convergence guarantees for the loss function. In contrast, we go one step further: by leveraging strong convexity, we establish finite-sample guarantees not just for the learning loss, but for convergence of the policy parameters. Our analysis is algorithm-agnostic and applies to any solver capable of optimizing the robust DPO loss. Additionally, our formulation uses KL and Wasserstein uncertainty sets, which are more standard in large-scale LLM alignment.
@@
Distributionally Robust Learning: Distributionally robust learning is a statistical learning framework designed to enhance model performance under distributional shifts between training and test data (Chen and Paschalidis,, 2018). It employs a minimax approach where an adversary maximizes the expected loss by shifting the test distribution within a specified uncertainty set, while the learner minimizes this adversarial loss. This approach using the $f$-divergence (Namkoong and Duchi,, 2016; Duchi and Namkoong,, 2021; Levy et al.,, 2020) and the Wasserstein metric (Mohajerin Esfahani and Kuhn,, 2018; Kuhn et al.,, 2019; Gao et al.,, 2022) have gained significant attention recently. Distributionally robust algorithms have been developed to address problems in supervised learning (Chen and Paschalidis,, 2018; Namkoong and Duchi,, 2016; Levy et al.,, 2020), multi-armed bandits (Si et al.,, 2020; Yang et al.,, 2023) and reinforcement learning (Panaganti et al.,, 2022; Zhou et al.,, 2024; Shi and Chi,, 2024; Yang et al.,, 2022).
@@ 3 Preliminaries
[…]
Wasserstein Distance: For a given set $\mathcal{Z}$, equipped with a metric $d$, the Wasserstein distance of order $p$ between two distributions $\mu,\nu\in\mathcal{P}(\mathcal{Z})$ is defined as (see Villani et al., (2009)):
@@
Kullback-Leibler Divergence: For any two probability distributions $\mathsf{P}$ and $\mathsf{Q}$ defined on $\mathcal{Z}$, the Kullback-Leibker (KL) divergence is defined as
$D_{\mathrm{KL}}(\mathsf{P}\;\|\;\mathsf{Q})=\sum_{z\in\mathcal{Z}}\mathsf{P}(z)\log(\mathsf{P}(z)/\mathsf{Q}(z)).$
@@
Step 2: Reward Modelling. In the second step, given any context $s\in\mathcal{S}$, two responses $a^{1},a^{2}\in\mathcal{A}$ are independently sampled from the behavior policy $\pi^{o}$, which is usually chosen as the SFT policy $\pi_{\mathrm{SFT}}$. Then, a (human) labeler provides a preference response between these responses. We assume that the preference responses are generated according to the Bradley-Terry (BT) model (Bradley and Terry,, 1952):
@@
$\mathcal{L}^{\mathrm{RLHF}}(r_{\phi};\mathcal{D})=-\mathbb{E}_{(s,a^{w},a^{l})\sim\mathcal{D}}[\log\sigma(r_{\phi}(s,a^{w})-r_{\phi}(s,a^{l}))]$
@@
Direct Preference Optimization (DPO): The DPO approach (Rafailov et al.,, 2023) leverages the fact that the unknown reward function can be expressed in terms of the optimal policy and the reference policy. Formally, given any reward function $r^{*}$, the optimal solution of Eq. 2 takes the form $\pi^{*}(a\mid s)=\frac{1}{Z^{*}(s)}\pi_{\mathrm{ref}}(a\mid s)\mathrm{exp}\left(r^{*}(s,a)/\beta\right)$, where $Z^{*}(s)$ denotes the partition (normalizing) function. Rearranging the above, we get $r^{*}(s,a)=\beta\log\frac{\pi^{*}(a\mid s)}{\pi_{\mathrm{ref}}(a\mid s)}+\beta\log Z^{*}(s)$ for all $(s,a)$. Substituting this into Eq. 1, the optimal RLHF policy $\pi^{*}$ satisfies the preference model:
@@ 4 Distributionally Robust DPO
In this section, we give the formulation of our Wasserstein DPO (WDPO) and Kullback-Leibler DPO (KLDPO).
Sampling Procedure: As described in Section 3, a prompt $s\in\mathcal{S}$ is drawn from an initial distribution $\mu^{o}$, and two responses $a^{1},a^{2}\sim_{\mathrm{i.i.d.}}\pi^{o}(\cdot\mid s)$ are sampled independently (with $\pi^{o}=\pi_{\mathrm{SFT}}$ in practice). Following Zhu et al., (2023), we define $y\in\{0,1\}$ to indicate preference: $y=1$ if $a^{1}\succ a^{2}\mid s$ and $y=0$ otherwise. The label $y$ is drawn from a Bernoulli distribution defined by the BT model $P^{*}$. The full data-generating distribution is given below.
@@ 4.1 Distributionally Robust DPO
From the DPO objective (Eq. 3), we define the pointwise DPO loss function as follows
@@
When we have a Wasserstein uncertainty set $\mathcal{P}_{\mathsf{W}_{p}}$, i.e., Eq. 4 equipped with the $p$-th order Wasserstein distance, we define the Wasserstein DPO (WDPO) loss as follows
@@
Similarly, given a Kullback-Leibler uncertainty set $\mathcal{P}_{\mathrm{KL}}(\rho;\mathsf{P}^{o})$, we define the KLDPO loss function as follows
@@ 5 Theoretical Analysis
This is a standard assumption in the theoretical analysis of the RL algorithms (Agarwal et al.,, 2021; Modi et al.,, 2020), RLHF (Zhu et al.,, 2023), and DPO (Nika et al.,, 2024; Chowdhury et al.,, 2024). Our analysis can be extended to the neural policy class where $\theta^{\top}\psi(s,a)$ is replaced $f_{\theta}(s,a)$, where $f_{\theta}$ is a neural network with twice differentiability and smoothness assumptions.
@@
We note that similar assumptions on data coverage under linear architecture models are standard in the offline RL literature (Agarwal et al.,, 2019; Wang et al.,, 2021; Jin et al.,, 2021). Implicitly, Assumption 2 imposes $\lambda\leq\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}^{o}})$, which means that the data-generating distribution $\mathsf{P}^{o}$ has good coverage.

### v3

@@ Abstract
A major challenge in aligning large language models (LLMs) with human preferences is the issue of distribution shift. LLM alignment algorithms rely on static preference datasets, assuming that they accurately represent real-world user preferences. However, user preferences vary significantly across geographical regions, demographics, linguistic patterns, and evolving cultural trends. This preference distribution shift leads to catastrophic alignment failures in many real-world applications. We address this problem using the principled framework of distributionally robust optimization, and develop two novel distributionally robust direct preference optimization (DPO) algorithms, namely, Wasserstein DPO (WDPO) and Kullback–Leibler DPO (KLDPO). We characterize the sample complexity of learning the optimal policy parameters for WDPO and KLDPO. Moreover, we propose scalable gradient descent-style learning algorithms by developing suitable approximations for the challenging minimax loss functions of WDPO and KLDPO. Our empirical experiments using benchmark data sets and LLMs demonstrate the superior performance of WDPO and KLDPO in substantially improving the alignment when there is a preference distribution shift.
@@ 1 Introduction
The alignment of large language models (LLMs) with human values and preferences is a central objective in machine learning, enabling these models to produce outputs that are useful, safe, and aligned with human intent. Since LLMs are trained on vast, diverse datasets using self-supervised learning, an additional alignment phase is often required to refine their behavior based on human feedback. A widely adopted approach for this is Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022), which involves training a reward model using human preference data and optimizing the LLM using reinforcement learning (RL) approaches, such as proximal policy optimization. More recently, Direct Preference Optimization (DPO) has emerged as an alternative that simplifies the alignment process by directly optimizing model parameters based on human preferences without requiring an explicit reward model. These alignment techniques have played a crucial role in improving the ability of LLMs to generate responses that adhere to human expectations and societal norms, leading to today’s powerful chat models (Achiam et al., 2023; Touvron et al., 2023).
@@
Despite the importance of the LLM alignment problem, RLHF and DPO remain fundamentally challenging and fragile, mainly due to three reasons. $(i)$ Diversity of human preferences: Standard RLHF/DPO approaches implicitly assume that human preferences can be accurately captured by a single reward function. In reality, human preferences are highly diverse, context-dependent, and distributional, making it infeasible to represent them with a one-size-fits-all optimization framework (Zhao et al., 2024; Durmus et al., 2024). Standard preference-learning methods tend to skew toward the preferences represented in the majority of training data, disproportionately penalizing minority opinions and reinforcing biases (Chakraborty et al., 2024). $(ii)$ Reward hacking: The quality of human preference feedback is inherently noisy, ambiguous, and inconsistent, as they are collected from human annotators who may lack domain expertise, exhibit labeling fatigue, or hold conflicting opinions (Zhang et al., 2025; Wu et al., 2025), which can often lead to misaligned preference estimation. This issue is exacerbated by reward hacking, where models learn undesirable shortcuts to maximize the estimated reward function, generating responses that appear aligned but deviate from genuine human intent (Amodei et al., 2016; Skalse et al., 2022; Eisenstein et al., 2024). $(iii)$ Distribution shift: Alignment algorithms use static preference datasets for training, collected under controlled conditions. However, the preferences of real-world users can often be out-of-distribution from that of the training data, depending on the geographical region, demography, linguistic patterns, and emerging social trends, among many others. A model aligned using a specific fixed dataset may fail catastrophically when deployed to users whose preference distribution does not match that of the training data (Casper et al., 2023; LeVine et al., 2023; Kirk et al., 2024).
@@
In this paper, we address the fragility of the LLM alignment using DPO, with a particular focus on the challenges arising from the preference distribution shift. DPO reduces the alignment problem to a supervised learning problem. It is known that the performance of supervised learning algorithms degrades significantly in the out-of-distribution setting (Taori et al., 2020; Koh et al., 2021), which is exacerbated due to the realistic distribution shift scenarios arising in the LLM deployment. Distributionally robust optimization/learning framework has been recently used to address the issue of distribution shift in various settings (Duchi and Namkoong, 2021; Kuhn et al., 2019; Chen et al., 2020). This framework considers an uncertainty set of data distributions around a nominal distribution (typically the training data distribution) and solves a minimax optimization problem to minimize the expected loss, where the expectation is taken with respect to the distribution in the uncertainty set that maximizes the loss. The distributionally robust learning approach has been successfully applied, with theoretical guarantees and scalable algorithms, in supervised learning (Chen and Paschalidis, 2018; Namkoong and Duchi, 2016; Levy et al., 2020), multi-armed bandits (Si et al., 2020; Yang et al., 2023), and reinforcement learning (Wang and Zou, 2022; Panaganti et al., 2022; Zhou et al., 2024; Xu et al., 2023). This motivates us to address the following questions:
@@
    Empirically, we show that standard DPO is sensitive to preference distribution shift, leading to degraded performance when training and evaluation rewards differ. In contrast, our robust variants, WDPO and KLDPO, consistently achieve superior performance across diverse alignment tasks. For example, we fine-tune LLaMA-3.2-1B/3B-Instruct and LLaMA-3.1-8B-Instruct models on prompts from the HelpSteer2 dataset (Wang et al., 2024b) using preferences generated by the ArmoRM reward model (Wang et al., 2024a), and evaluate them on distinct reward objectives from the OpenLLM Leaderboard (Fourrier et al., 2024).
@@ 2 Related Work
Robust RLHF: Bai et al. (2022) proposed to adjust weights on the combination of loss functions based on different topics (harmless vs. helpful) for robust reward learning. Chakraborty et al. (2024) proposed to learn multiple reward functions for different sub-populations through an expectation-maximization approach, and a robust policy based on these rewards via a max-min optimization, which is different from our distributional robust learning approach. Padmakumar et al. (2024) augmented the existing binary preference datasets with synthetic preference judgments to estimate the diversity of user preferences. Yan et al. (2024) proposed a Bayesian reward model ensemble to quantify the uncertainty of reward estimation and used it to reduce reward overoptimization. Bukharin et al. (2024) proposed a robust RLHF approach for addressing the preference data corruption problem.
[…]
@@
Distributionally Robust Learning: Distributionally robust learning is a statistical learning framework designed to enhance model performance under distributional shifts between training and test data (Chen and Paschalidis, 2018). It employs a minimax approach where an adversary maximizes the expected loss by shifting the test distribution within a specified uncertainty set, while the learner minimizes this adversarial loss. This approach using the $f$-divergence (Namkoong and Duchi, 2016; Duchi and Namkoong, 2021; Levy et al., 2020) and the Wasserstein metric (Mohajerin Esfahani and Kuhn, 2018; Kuhn et al., 2019; Gao et al., 2022) have gained significant attention recently. Distributionally robust algorithms have been developed to address problems in supervised learning (Chen and Paschalidis, 2018; Namkoong and Duchi, 2016; Levy et al., 2020), imitation learning (Bashiri et al., 2021; Panaganti et al., 2023), multi-armed bandits (Si et al., 2020; Yang et al., 2023), and reinforcement learning (Panaganti et al., 2022; Zhou et al., 2024; Shi and Chi, 2024; Yang et al., 2022; Panaganti et al., 2025).
@@ 3 Preliminaries
Notations: We use calligraphic letters for sets, e.g., $\mathcal{S}$. $\lVert\cdot\rVert$ denotes the Euclidean norm. When $\Sigma$ is a positive semi-definite matrix, we write $\lVert x\rVert_{\Sigma}=\sqrt{x^{\top}\Sigma x}$ as a semi-norm of $x$. For any measure $\mathsf{P}$, we use $\mathsf{P}_{n}$ to denote the empirical distribution constructed using $n$ i.i.d. samples, $x_{1},\dots,x_{n}$, from $\mathsf{P}$, i.e., $\mathsf{P}_{n}=(1/n)\sum_{i=1}^{n}\delta_{x_{i}}$, where $\delta_{x}$ is the Dirac measure. We use $\sigma$ to denote the sigmoid (standard logistic) function. We use $l(z;\theta)$ and $l_{z}(\theta)$ to denote the loss incurred by sample $z$ with policy parameter $\theta$. For any set $\mathcal{Z}$, $\mathcal{P}(\mathcal{Z})$ is the set of all Borel measures over $\mathcal{Z}$. For any positive semi-definite matrix $\Sigma$, $\operatorname{\lambda_{\mathrm{min}}}(\Sigma)$ and $\operatorname{\lambda_{\mathrm{max}}}(\Sigma)$ denote its smallest and largest eigenvalues.
Wasserstein Distance: For a given set $\mathcal{Z}$, equipped with a metric $d$, the Wasserstein distance of order $p$ between two distributions $\mu,\nu\in\mathcal{P}(\mathcal{Z})$ is defined as (see Villani et al. (2009)):
@@
Kullback-Leibler Divergence: For any two probability distributions $\mathsf{P}$ and $\mathsf{Q}$ defined on $\mathcal{Z}$, the Kullback-Leibker (KL) divergence is defined as $D_{\mathrm{KL}}(\mathsf{P}\;\|\;\mathsf{Q})=\sum_{z\in\mathcal{Z}}\mathsf{P}(z)\log(\mathsf{P}(z)/\mathsf{Q}(z)).$
@@
Step 2: Reward Modelling. In the second step, given any context $s\in\mathcal{S}$, two responses $a^{1},a^{2}\in\mathcal{A}$ are independently sampled from the behavior policy $\pi^{o}$ (typically the SFT policy $\pi_{\mathrm{SFT}}$). Then, a (human) labeler provides a preference response between these responses. We assume that the preference responses are generated according to the Bradley-Terry (BT) model (Bradley and Terry, 1952):
@@
$\mathcal{L}^{\mathrm{RLHF}}(r_{\phi};\mathcal{D})=-\mathbb{E}_{(s,a^{w},a^{l})\sim\mathcal{D}}[\log\sigma(r_{\phi}(s,a^{w})-r_{\phi}(s,a^{l}))].$
@@
Direct Preference Optimization (DPO): The DPO approach (Rafailov et al., 2023) leverages the fact that the unknown reward function can be expressed in terms of the optimal policy and the reference policy. Formally, given any reward function $r^{*}$, the optimal solution of Eq.˜2 takes the form $\pi^{*}(a\mid s)=\frac{1}{Z^{*}(s)}\pi_{\mathrm{ref}}(a\mid s)\mathrm{exp}\left(r^{*}(s,a)/\beta\right)$, where $Z^{*}(s)$ denotes the partition (normalizing) function. Rearranging the above, we get $r^{*}(s,a)=\beta\log\frac{\pi^{*}(a\mid s)}{\pi_{\mathrm{ref}}(a\mid s)}+\beta\log Z^{*}(s)$ for all $(s,a)$. Substituting this into Eq.˜1, the optimal RLHF policy $\pi^{*}$ satisfies the preference model:
@@ 4 Distributionally Robust DPO
In this section, we formulate our Wasserstein DPO (WDPO) and Kullback-Leibler DPO (KLDPO).
Sampling Procedure: As described in Section˜3, a prompt $s\in\mathcal{S}$ is drawn from an initial distribution $\mu^{o}$, and two responses $a^{1},a^{2}\sim_{\mathrm{i.i.d.}}\pi^{o}(\cdot\mid s)$ are sampled independently (with $\pi^{o}=\pi_{\mathrm{SFT}}$ in practice). Following Zhu et al. (2023), we define $y\in\{0,1\}$ to indicate preference: $y=1$ if $a^{1}\succ a^{2}\mid s$ and $y=0$ otherwise. The label $y$ is drawn from a Bernoulli distribution defined by the BT model $P^{*}$. The full data-generating distribution is given below.
@@ 4.1 Distributionally Robust DPO
From the DPO objective (Eq.˜3), we define the pointwise DPO loss function as follows
@@
When we have a Wasserstein uncertainty set $\mathcal{P}_{\mathsf{W}_{p}}$, i.e., Eq.˜4 equipped with the $p$-th order Wasserstein distance, we define the Wasserstein DPO (WDPO) loss as follows
@@
Similarly, given a Kullback-Leibler uncertainty set $\mathcal{P}_{\mathrm{KL}}(\rho;\mathsf{P}^{o})$, we define the KLDPO loss as follows
@@ 5 Theoretical Analysis
This is a standard assumption in the theoretical analysis of the RL algorithms (Agarwal et al., 2021; Modi et al., 2020), RLHF (Zhu et al., 2023), and DPO (Nika et al., 2024; Chowdhury et al., 2024). Our analysis can be extended to the neural policy class where $\theta^{\top}\psi(s,a)$ is replaced $f_{\theta}(s,a)$, where $f_{\theta}$ is a neural network with twice differentiability and smoothness assumptions.
@@
We note that similar assumptions on data coverage under linear architecture models are standard in the offline RL literature (Agarwal et al., 2019; Wang et al., 2021; Jin et al., 2021). Implicitly, Assumption˜2 imposes $\lambda\leq\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}^{o}})$, which means that the data-generating distribution $\mathsf{P}^{o}$ has good coverage.